l. An experiment has four equiprobable outcomes, S-{ 1 ,2, 3, 4). We define three events: A-{ I , 2), В-{ 2, 3 } and C-{ l , 3). Are A, B and C independent? Justify your answer.
Answer #1

Now,

P(A) = P(B) = P(C) =-=

and

A cap B = {1,2} cap {2,3} = {2} that implies P(Acap B) = rac{1}{4}

Here,

P(AnB)P(A) P(B)

So, A and B are independent.

Similarly,

A cap C = {1,2} cap {1,3} = {1} that implies P(Acap C) = rac{1}{4}

Here,

P(Acap C) = rac{1}{4} = P(A)*P(C)

So, A and C are independent.

and also

B cap C = {2,3} cap {1,3} = {3} that implies P(Acap C) = rac{1}{4}

Here,

P(Bcap C) = rac{1}{4} = P(B)*P(C)

So, B and C are also independent.

Hence, A, B and C are mutually independent.